# CholecInstanceSeg: A Tool Instance Segmentation Dataset for Laparoscopic Surgery

### v1

@@ Abstract
In laparoscopic and robotic surgery, precise tool instance segmentation is an essential technology for advanced computer-assisted interventions. Although publicly available procedures of routine surgeries exist, they often lack comprehensive annotations for tool instance segmentation. Additionally, the majority of standard datasets for tool segmentation are derived from porcine(pig) surgeries. To address this gap, we introduce CholecInstanceSeg, the largest open-access tool instance segmentation dataset to date. Derived from the existing CholecT50 and Cholec80 datasets, CholecInstanceSeg provides novel annotations for laparoscopic cholecystectomy procedures in patients. Our dataset comprises 41.9k annotated frames extracted from 85 clinical procedures and 64.4k tool instances, each labelled with semantic masks and instance IDs. To ensure the reliability of our annotations, we perform extensive quality control, conduct label agreement statistics, and benchmark the segmentation results with various instance segmentation baselines. CholecInstanceSeg aims to advance the field by offering a comprehensive and high-quality open-access dataset for the development and evaluation of tool instance segmentation algorithms. CholecInstanceSeg is publicly available here
@@
To overcome these challenges, we introduce the CholecInstanceSeg open-access dataset, in which we meticulously annotate 41,933 frames extracted from 85 laparoscopic cholecystectomy procedures from the existing Cholec80 [12] and CholecT50 [13] datasets. To the best of our knowledge, CholecInstanceSeg represents the most extensive surgical tool instance segmentation dataset. Table 1 compares popular datasets used in tool segmentation research and CholecInstanceSeg. We acknowledge a concurrent effort, the PhaKIR dataset, which is being developed for the upcoming MICCAI2024 PhaKIR-challenge. The PhaKIR dataset is expected to include 15 videos and approximately 30,000 instance annotations. However, as this dataset is only partially released and not yet available for research, we have not included it in our comparisons.
@@
We provide technical validations of our dataset, demonstrating our comprehensive quality control procedures and showcasing CholecInstanceSeg’s capability for training instance segmentation models. CholecInstanceSeg is ideally suited for training models on instance and semantic segmentation for laparoscopic surgeries. Additionally, it can be effectively integrated with other datasets that utilize surgical procedure videos from Cholec80 and CholecT50. These datasets, derived from Cholec80 and CholecT50 share frames with CholecInstanceSeg but are annotated for different tasks, such as phase recognition (Cholec80 [12]) instrument action and target detection (CholecT50 [13]), and Critical View of Safety (CVS) assessment (Cholec80-CVS [14]).
@@ Methods
The CholecInstanceSeg dataset is curated using frames extracted from three existing laparoscopic cholecystectomy datasets: Cholec80 [12], CholecT50 [13], and CholecSeg8k [11], containing a total of 85 image sequences. These datasets are publicly available subsets of the in-house Cholec120 dataset [15] from the CAMMA research group at the University of Strasbourg. The interrelation between these datasets is illustrated in Figure 1.
@@
CholecInstanceSeg, which contains 41.9k frames from 85 unique image sequences, can be partitioned into four distinct sections based on the data source: Instance-CholecSeg8k, Instance-CholecT50-full, Instance-CholecT50-sparse, and Instance-Cholec80-sparse.
@@
    Instance-CholecSeg8k: For the 8,080 frames over 17 sequences provided in CholecSeg8k, we utilized all the images in this dataset.
@@
    Instance-CholecT50-full: For 15 sequences within CholecT50(CholecT50-full), we utilized all frames provided by CholecT50.
@@
    Instance-CholecT50-sparse: For the remaining 35 sequences in CholecT50(CholecT50-sparse), we sampled these sequences by selecting one frame out of every 30 frames ($\frac{1}{30}$ fps), ensuring a minimum of 50 frames in each sequence.
@@
    Instance-Cholec80-sparse: We extracted 28 videos from Cholec80, which are not included in CholecT50 and CholecSeg8k, and sampled these videos at $\frac{1}{30}$ fps, ensuring a minimum of 50 frames in each sequence.
@@
The inclusion of the snare category deviates from the standard six categories (grasper, bipolar, hook, clipper, scissors, irrigator) defined by Cholec80 and CholecT50. We determined that the snare’s distinct visual characteristics and unique functionality warranted its classification as a separate tool category. A visual representation of these tool categories is shown in Figure 3
Figure 3: Visual representation of the tool categories used in the CholecInstanceSeg dataset. Each image shows a different tool: Grasper, Snare, Irrigator, Clipper, Scissors, Bipolar, and Hook.
@@
    A clear visual representation of this unique situation with the instrument port can be seen in Figure 4
Figure 4: A visual representation of the unique annotation situation with the instrument port. On the left, no annotation is made when the instrument is not seen extending from the port. On the right, the instrument and the port are annotated as a single entity when the instrument extends from the port.
@@
Figure 5: Visual representation of challenging annotation scenarios. Each image highlights a specific hard case: motion blur, smoke, soft tissue occlusion, tissue attachment, saturated lighting, instrument at the edge, reflection, instrument in fluid, low lighting, dirty lens, camera in port, and instrument far from the camera.
@@
  • Annotation of Hard Cases: CholecInstanceSeg captures routine surgical procedures, numerous challenging scenarios were encountered. We accounted for twelve specific situations: motion blur, presence of smoke, occlusion by soft transparent tissues or blood, tissues attached to instruments, instruments at the edge of the frame, instruments far from the camera, light reflection, low lighting conditions, bright lighting conditions, lens dirtiness, camera in camera port, and instruments in liquid. Specialized instructions and examples were provided to annotators, including checking neighbouring frames for clarification, utilizing the expected shape of the instrument when in doubt, annotating as much of the instrument as possible, and annotating even under low visibility conditions. A visual representation of these hard cases can be seen in Figure 5. More detailed information about specialized instructions given to annotators for hard cases is added as supplementary material for this manuscript.
@@
    Annotation review and correction: Human annotators reviewed the model-generated annotations and corrected any errors or inaccuracies, ensuring the quality and accuracy of the annotations.
@@
The CholecInstanceSeg dataset is publicly available and can be accessed on synapse. This section presents data analysis to show some of the properties of CholecInstanceSeg and provides information about the recommended splits, the dataset’s structure and content.
@@
The final split for the dataset is as follows: the training set contains 55 sequences (26,830 frames), the testing set includes 23 sequences (11,299 frames), and the validation set comprises 17 sequences (3,804 frames).
@@
File Format
CholecInstanceSeg is provided as a single zip file, organized into two sub-directories: train and val directories. The test directory is currently withheld for benchmarking purposes and potential competition use. The testing dataset is available on request.
Each sub-directory contains a directory for each video sequence, which in turn contains two directories: one for images (img_dir) and another for annotations (ann_dir). Images are in PNG format, while annotation files are in JSON format.
Annotations follow the LabelMe format. Each annotated shape is listed under the "shapes" key, with a "label" denoting the class, "points" specifying the (x,y) coordinates of the polygon, and a "group_id" for the instance ID.
CholecInstanceSeg shares frames with Cholec80, Cholec80-CVS, CholecSeg8k, and CholecT50 datasets. To maintain consistency, video names are preserved and frame numbers retain the same IDs as the source datasets, with prefixes added to indicate the sequence number and dataset partition. frame numbers (e.g., VID01, VID70) are preserved to enable researchers to combine annotations between our dataset and these datasets.
Most frames in CholecInstanceSeg are 854x480 pixels, except for three sequences in Instance-CholecT50-full, which are 1920x1080 pixels(same as the source dataset). We provide scripts to resize these images and corresponding annotations to 854x480 pixels if needed. A schematic diagram of the folder structure and a sample JSON annotation is shown in Figure 10.
Figure 10: (a) Dataset directory structure for CholecInstanceSeg. Each split (e.g. train) contains subdirectories for each sequence, with further subdirectories for images ($img\_dir$) and annotations ($ann\_dir$). (b) Sample JSON annotation file showing the structure of the annotation data, including class labels ("$label$"), polygon information ("$points$") and instance IDs ("$group\_id$").
Technical Validation
We provide an in-depth technical validation of the CholecInstanceSeg dataset, demonstrating its reliability and effectiveness for use in research and development of instance segmentation models. We discuss the validation through three main aspects: training of baseline models, label agreement statistics, and quality control measures.
Training and Evaluation of Baseline Models
Models
@@
To evaluate our models we used the standard COCO mean average precision metric (COCO mAP), and a metric which we call the COCO sequence mean average precision (COCO smAP).
From the dataset analysis, we could tell that our dataset has a class imbalance which is captured by the mean average precision, however, it also has an imbalance among sequences, such that some sequences have a lot of tool instances and images while others do not. Hence we wanted to also have a metric that captured how well models are performing on sequences. It involves a slight change to the COCO mean average precision metric. COCO smAP generates the average precision per class over each of the sequences separately and then averages this value for the whole dataset. This metric enables the penalization of models that do not generalize to different sequences and scenes.
The sequence Average Precision ($sAP_{j}$) for a class $j$ (COCO smAP can be written as
@@
Here, $s$ represents the current sequence, $N$ represents the number of sequences evaluated, and $p_{seq=i}(j,r)$ represents the precision at recall $r$ for class $j$ over a sequence $s$.
We utilize the official COCO implementation present here for calculating the COCO mAP metric. To determine the COCO smAP metric, we first compute the COCO mAP for each sequence separately and average the results for all sequences.
@@ Results
The performance results, presented in Table 3, show the effectiveness of the dataset in training robust instance segmentation models. A figure showing the qualitative results of both models against groundtruth for an image from the validation dataset can be seen at Figure 11. A video of more results from both models on the validation dataset can be found here
Figure 11: Visual representation of the performance of Mask R-CNN, Mask2former and the expected ground truth.
Model val mAP val smAP
Mask R-CNN 0.470 0.437
Mask2Former 0.682 0.655
Table 3: Performance of baseline instance segmentation models on CholecInstanceSeg.
@@
Usage Notes
CholecInstanceSeg is released under a Creative Commons Attribution-NonCommercial ShareAlike license (CC BY-NC-SA). Any use or mention of this dataset must include a citation to this paper. If the dataset is utilized in the creation of new works, they should also include citations to this paper.
This licensing choice aligns with the release license of the Cholec80 dataset, from which our frames are derived, ensuring that the original license is retained and respected.
@@
The relevant code for the use of this dataset is found here. This repository includes scripts that provide information about class IDs, sequences, and splits. It also contains notebooks for converting between dataset partitions and splits, utilities for converting to the COCO format [25], visualization tools, and additional resources. The official implementation of COCO mAP and panoptic quality metric were used and they can be found on the COCO dataset GitHub page.
@@ Acknowledgements
0.A. wrote the original draft, was the primary annotator, and conducted the experiments and tasks that led to this dataset, K.Z.T. was the secondary annotator for this project, assisted with quality control, and the creation of the annotation guide. Z.Z., C.B., M.S., T.V. were part of the annotation team and mediated ambiguous labels. M.S. and T.V. conceived and supervised this project.

### v2

@@ Abstract
In laparoscopic and robotic surgery, precise tool instance segmentation is an essential technology for advanced computer-assisted interventions. Although publicly available procedures of routine surgeries exist, they often lack comprehensive annotations for tool instance segmentation. Additionally, the majority of standard datasets for tool segmentation are derived from porcine(pig) surgeries. To address this gap, we introduce CholecInstanceSeg, the largest open-access tool instance segmentation dataset to date. Derived from the existing CholecT50 and Cholec80 datasets, CholecInstanceSeg provides novel annotations for laparoscopic cholecystectomy procedures in patients. Our dataset comprises 41.9k annotated frames extracted from 85 clinical procedures and 64.4k tool instances, each labelled with semantic masks and instance IDs. To ensure the reliability of our annotations, we perform extensive quality control, conduct label agreement statistics, and benchmark the segmentation results with various instance segmentation baselines. CholecInstanceSeg aims to advance the field by offering a comprehensive and high-quality open-access dataset for the development and evaluation of tool instance segmentation algorithms.
@@
To overcome these challenges, we introduce the CholecInstanceSeg open-access dataset, in which we meticulously annotate 41,933 frames from the existing Cholec80 [12] and CholecT50 [13] datasets with instance segmentation labels and class labels for 7 instrument classes: Grasper, Bipolar, Hook, Clipper, Scissors, Irrigator, and Snare. CholecInstanceSeg contains frames from 85 videos of Laparoscopic cholecystectomy - a routine surgical procedure for gallbladder removal, commonly performed to treat gallstones or cholecystitis with over 300,000 surgeries performed annually in the United States [14]. To the best of our knowledge, CholecInstanceSeg represents the most extensive surgical tool instance segmentation dataset. Table 1 compares popular datasets used in tool segmentation research and CholecInstanceSeg. We acknowledge a concurrent effort, the PhaKIR dataset, which is being developed for the upcoming MICCAI2024 PhaKIR-challenge. The PhaKIR dataset is expected to include 15 videos and approximately 30,000 instance annotations. However, as this dataset is only partially released and not yet available for research, we have not included it in our comparisons.
@@
Furthermore, we provide technical validations of our dataset, demonstrating our comprehensive quality control procedures and showcasing CholecInstanceSeg’s capability for training instance segmentation models.
CholecInstanceSeg is ideally suited for training models on instance and semantic segmentation for laparoscopic surgeries. Additionally, it can be effectively integrated with other datasets that utilize surgical procedure videos from Cholec80 and CholecT50. These datasets, derived from Cholec80 and CholecT50 share frames with CholecInstanceSeg but are annotated for different tasks, such as phase recognition (Cholec80 [12]) instrument action and target detection (CholecT50 [13]), and Critical View of Safety (CVS) assessment (Cholec80-CVS [15]).
@@ Methods
The CholecInstanceSeg dataset is curated using frames extracted from three existing laparoscopic cholecystectomy datasets: Cholec80 [12], CholecT50 [13], and CholecSeg8k [11], containing a total of 85 image sequences. These datasets are publicly available subsets of the in-house Cholec120 dataset [16] from the CAMMA research group which was recorded at the University Hospital of Strasbourg [12, 13] . The interrelation between these datasets is illustrated in Figure 1.
@@
The frames from Cholec80 and CholecT50 used in CholecInstanceSeg are primarily at a resolution of 854x480 pixels, consistent with the original Cholec80 and CholecT50 datasets. However, three sequences from CholecT50 are provided at a higher resolution of 1920x1080 pixels. CholecT50 contains pre-extracted image sequences and we used these as is, while frames from Cholec80 were extracted at 1 fps using the FFMPEG library.
@@
CholecInstanceSeg, which contains 41.9k frames from 85 unique image sequences, can be partitioned into four distinct sections based on the data source: Instance-CholecSeg8k, Instance-CholecT50-full, Instance-CholecT50-sparse, and Instance-Cholec80-sparse. Our process for selecting the frames and sequences for each partition is as follows:
@@
    Instance-CholecSeg8k: We selected all the 8,080 frames from all the 17 sequences provided in CholecSeg8k, we utilized all the images in CholecSeg8k.
@@
    Instance-CholecT50-full: We randomly selected 15 sequences within CholecT50 (CholecT50-full), we utilized all frames provided by CholecT50 for these 15 sequences - 28,317 frames.
@@
    Instance-CholecT50-sparse: For the remaining 35(50-15) sequences in CholecT50(CholecT50-sparse), we sampled these sequences by selecting one frame out of every 30 frames ($\frac{1}{30}$ fps), ensuring a minimum of 50 frames in each sequence - 2,681 frames .
@@
    Instance-Cholec80-sparse: After selecting sequences for Instance-CholecSeg8k, Instance-CholecT50 (both full and sparse), there remained 28 videos from Cholec80, which are not included in CholecT50 and CholecSeg8k, and these videos were sampled at $\frac{1}{30}$ fps, ensuring a minimum of 50 frames in each sequence - 2,855 frames .
@@
The differences in frame rates for each dataset partition reflect both the design of the parent datasets – Cholec80 provides videos at 25fps which we then extract at 1fps and CholecT50 provides image sequences at 1fps directly – and practical considerations for annotating instance segmentation on a large number of frames. Annotating every sequence densely at 1fps is prohibitively expensive. Thus, we adopted a mixed approach, converting existing semantic segmentation to instance segmentation, annotating some sequences fully (using all frames available from the parent datasets), and others sparsely (sampling frames at reduced frame rates of $\frac{1}{30}$ fps). This approach allows us to balance annotation density and dataset diversity, ensuring coverage of both densely annotated sequences and a wider range of surgical scenes.
Figure 7 presents the exact sequences in each partition, represented as VID-XX, where XX is the sequence number.
@@
The inclusion of the snare category deviates from the standard six categories (grasper, bipolar, hook, clipper, scissors, irrigator) defined by Cholec80 and CholecT50. We determined that the snare’s distinct visual characteristics and unique functionality warranted its classification as a separate tool category.
@@
  • Annotation of Hard Cases: CholecInstanceSeg captures routine surgical procedures, numerous challenging scenarios were encountered. We accounted for twelve specific situations: motion blur, presence of smoke, occlusion by soft transparent tissues or blood, tissues attached to instruments, instruments at the edge of the frame, instruments far from the camera, light reflection, low lighting conditions, bright lighting conditions, lens dirtiness, camera in camera port, and instruments in liquid. Specialized instructions and examples were provided to annotators, including checking neighbouring frames for clarification, utilizing the expected shape of the instrument when in doubt, annotating as much of the instrument as possible, and annotating even under low visibility conditions.
@@
    Annotation review and correction: Human annotators reviewed the model-generated annotations and corrected any errors or inaccuracies, ensuring the quality and accuracy of the annotations. Minor errors were corrected with the capabilities of the LabelMe labeling tool, while major errors requiring new annotations utilized the point-prompt method of the SAM model.
@@
The CholecInstanceSeg dataset [21] is publicly available. Cholec80 and CholecT50 video frames used to generate the dataset can be obtained from the CAMMA research group website(http://camma.u-strasbg.fr/datasets). CholecSeg8k can be obtained at (https://www.kaggle.com/datasets/newslab/cholecseg8k).
The dataset is provided as a single archive, $CholecInstanceSeg.zip$, which contains the instance segmentation annotation files in $cholecinstanceseg.zip$ and a metadata file $cholecinstanceseg\_metadata.csv$ that provides metadata information about each annotation file and can be used to link each file to its source dataset. The annotation files are structured into two main directories: train and val, each containing individual video sequence folders. Each sequence folder includes an $ann\_dir$ subdirectory that stores instance segmentation annotations in JSON format. The test directory is withheld for benchmarking purposes and potential competition use, but it is available upon request. A schematic representation of the dataset folder structure and a sample JSON annotation file are shown in Figure 4
Annotations are formatted following the LabelMe convention. Each shape is listed under the $shapes$ key, with a $label$ field specifying the instrument class, $points$ containing the (x, y) coordinates of the polygon defining the object’s boundary, $group\_id$ indicating the instance ID, $image\_Height$ for the height of image annotated and $image\_Width$ for its width.
Figure 4: (a) Dataset directory structure for CholecInstanceSeg. Each split (e.g. train) contains subdirectories for each sequence, with further subdirectories for annotations ($ann\_dir$). (b) Sample JSON annotation file showing the structure of the annotation data, including class labels ("$label$"), polygon information ("$points$") and instance IDs ("$group\_id$").
Technical Validation
We provide an in-depth technical validation of the CholecInstanceSeg dataset, demonstrating its reliability and effectiveness for use in research and development of instance segmentation models. We discuss the validation through three main aspects: Data Analysis, training of baseline models, label agreement statistics, and quality control measures.
@@
The final split for the dataset is as follows: the training set contains 55 sequences (26,830 frames), the testing set includes 22 sequences ( 10,819 frames), and the validation set comprises 18 sequences (4,284 frames ).
@@
To evaluate our models we used the standard COCO mean average precision metric (COCO mAP r@[0.50:0.95]). .
From the dataset analysis, we could tell that our dataset has a class imbalance which is captured by the mean average precision, however, it also has an imbalance among sequences, such that some sequences have a lot of tool instances and images while others do not. Hence we wanted to also have a metric that captured how well models are performing on sequences.
COCO smAP involves a slight modification of the COCO mAP metric. The sequence average precision(sAP) for a class $j$ over the whole dataset ($sAP_{j}$) in COCO smAP can be written as
@@
where $s$ represents the current sequence, $N$ represents the number of sequences evaluated, and $p_{seq=i}(j,r)$ represents the precision at recall $r$ for class $j$ over a sequence $s$. Essentially $sAP_{j}$ is calculated by taking the mean of the COCO average precision for that class for each sequence.
To measure variability in performance across sequences, we introduced the standard deviation of sequence average precision ($\sigma_{sAP_{j}}$).
$\sigma_{sAP_{j}}=\sqrt{\frac{1}{N}\sum_{s=0}^{N}\left(\left(\int_{0}^{1}p_{seq=s}(j,r)\right)\,dr-sAP_{j}\right)^{2}}$
where $s$ represents the current sequence, $N$ represents the number of sequences evaluated, and $p_{seq=i}(j,r)$ represents the precision at recall $r$ for class $j$ over a sequence $s$. $\sigma_{sAP_{j}}$ is calculated by taking the standard deviation of the COCO average precision for that class for each sequence.
Finally, the COCO smAP is calculated as a mean of the sequence average precisions across all classes
$smAP=\frac{1}{C}\sum_{j=1}^{C}sAP_{j}$
We utilize the official COCO implementation present here for calculating the COCO mAP and the metric COCO smAP metric.
@@ Results
The performance results, presented in Table 3 and Table 4, validate the dataset’s effectiveness in training robust instance segmentation models.
Notably, the hook class achieved the highest segmentation performance across both models, likely due to its visually distinct features and higher frequency in the dataset. In contrast, the scissor class exhibited the lowest performance, reflecting its smaller representation and potential challenges in detecting this class. Interestingly, the scissor class demonstrated a higher sequence mAP (smAP) but with significant variability (high standard deviation), suggesting that model performance on this class varies considerably across different sequences. Increasing the number and diversity of annotated sequences could improve dataset variability and enhance its applicability to a broader range of surgical scenarios, ensuring fewer seen instruments are better recognized by trained models.
Table 3: Performance of baseline instance segmentation models on CholecInstanceSeg for the mean average precision metric (mAP) providing the average precision for each class. GR - Grasper, HO - Hook, IR - irrigator, CL - clipper, BI - bipolar, SC - scissors, SN - snare.
Method GR HO IR CL BI SC SN mAP
Mask-RCNN [22] 37.25 61.84 43.04 58.73 28.42 20.91 - 41.70
Mask2former [23] 61.10 82.16 68.38 81.30 48.02 27.89 - 61.47
Table 4: Performance of baseline instance segmentation models on CholecInstanceSeg for the sequence mean average precision (smAP) providing the sequence average precision for each class j ($sAP_{j}$). GR - Grasper, HO - Hook, IR - Irrigator, CL - Clipper, BI - Bipolar, SC - Scissors, SN - Snare. The Standard deviation of the sequence average precision $\sigma_{sAP_{j}}$ is provided using the ($\pm$) symbol.
Method GR HO IR CL BI SC SN smAP
Mask-RCNN $41.65\pm 12.74$ $64.06\pm 9.20$ $52.50\pm 20.09$ $64.85\pm 16.20$ $37.62\pm 20.74$ $40.38\pm 23.54$ $50.18$
Mask2former $66.46\pm 8.52$ $88.90\pm 7.58$ $74.51\pm 20.35$ $84.69\pm 16.88$ $65.20\pm 25.36$ $69.76\pm 33.20$ $74.92$
@@
The relevant code for the use of this dataset is found here. This repository includes scripts that provide information about class IDs, sequences, and splits. It also contains notebooks for converting between dataset partitions and splits, utilities for converting to the COCO format [27], visualization tools, retrieving the input images, and additional resources. The official implementation of COCO mAP and panoptic quality metric were used and they can be found on the COCO dataset GitHub page.
@@ Acknowledgements
O.A. wrote the original draft, was the primary annotator, and conducted the experiments and tasks that led to this dataset, K.Z.T. was the secondary annotator for this project, assisted with quality control, and the creation of the annotation guide. Z.Z., C.B., M.S., T.V. were part of the annotation team and mediated ambiguous labels. M.S. and T.V. conceived and supervised this project.